# A distance function for stochastic matrices

### v1

@@ V Conclusion
However these distances explicitly require the initial distributions of a chain and can be extremely sensitive to small changes. We therefore considered a distance on the space of stochastic matrices, which can give a more robust measure of similarity. Moreover, it is valid for any pair of stochastic matrices, is computable and through Eq. 21 in the case of ergodic chains the asymptotic step distance is the same as considering distances on the sequence space. We were also able to find compact analytical results for very broad classes of stochastic matrices, namely matrix products which are diagonalisable or primitive for sequences spaces, and ergodoic chains for the stochastic matrix distance. Using underlying local structures (i.e. state to state jumps) to understand Markov model similarity this sense is more akin to comparing Hamiltonians or Lagrangians in physics, rather than looking at the whole possible trace space dependent on initial distributions.
@@
Despite the number of results, there is still much more worthwhile future work. For sequence spaces, finding the most general conditions for the matrix $\boldsymbol{R}$ to give a closed form solution for $d^{\omega}(M_{1},M_{2})$ would be useful. We expect the omission of eigenvalues of $-1$ from the spectrum of $\boldsymbol{R}$ can be dropped, as we know the Bhattacharyya angle is always convergent. Also the exploration of mixing times and the application of the Bhattacharyya angle to general mixing time bounds has only received an elementary review here. Tighter bounds are surely possible. Applications of the Bhattacharyya angle, and its stochastic matrix counterpart, also seem readily available to machine learning in cases where the problem requires the evaluation of a model which outputs categorical probabilities. As both distances are both analytically and numerically approachable, it would be pleasant to see it incorporated in to computatiaonal frameworks, such as for gradient boosting [54].
Acknowledgment
The authors would like to thank Sam Power, George Deligiannidis, Christopher Yau for their constructive and supportive comments for this paper. ARL would like to thank Sara Oriana Gomes Tavares for their very elegant argument on the convergence of the Bhattacharyya angle, which much improved upon his own.
Funding
ARL acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme in Health Data Science (Grant Ref: 218529/Z/19/Z).

### v2

@@ V Conclusion
However these distances explicitly require the initial distributions of a chain and can be extremely sensitive to small changes. We therefore considered a distance on the space of stochastic matrices, which can give a more robust measure of similarity. Moreover, it is valid for any pair of stochastic matrices, is computable and through Eq. 21 in the case of ergodic chains the asymptotic step distance is the same as considering distances on the sequence space. We were also able to find compact analytical results for very broad classes of stochastic matrices, namely matrix products which are diagonalisable or primitive for sequences spaces, and ergodoic chains for the stochastic matrix distance. Using underlying local structures (i.e. state to state jumps) to understand Markov chain similarity in this sense is more akin to comparing Hamiltonians or Lagrangians in physics, rather than looking at the whole possible trace space dependent on initial distributions.
@@
Despite the number of results, there is still much more worthwhile future work. For sequence spaces, finding the most general conditions for the matrix $\bm{R}$ to give a closed form solution for $d^{\omega}(M_{1},M_{2})$ would be useful. We expect the omission of eigenvalues of $-1$ from the spectrum of $\bm{R}$ can be dropped, as we know the Bhattacharyya angle is always convergent. Also the exploration of mixing times and the application of the Bhattacharyya angle to general mixing time bounds has only received an elementary review here. Tighter bounds are surely possible. Applications of the Bhattacharyya angle, and its stochastic matrix counterpart, also seem readily available to machine learning in cases where the problem requires the evaluation of a model which outputs categorical probabilities. As both distances are both analytically and numerically approachable, it would be pleasant to see it incorporated in to computational frameworks, such as for gradient boosting [56].
Acknowledgements.
The authors would like to thank Sam Power, George Deligiannidis, Christopher Yau for their constructive and supportive comments for this paper. ARL would like to thank Sara Oriana Gomes Tavares for their very elegant argument on the convergence of the Bhattacharyya angle, which much improved upon his own. ARL acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme in Health Data Science (Grant Ref: 218529/Z/19/Z). P. Tiňo was supported by the EPSRC Prosperity Partnerships grant ARCANE, EP/X025454/1.
Appendix A Proof of theorem 2
Proof.
Using the product structure of stochastic matrices, we can decompose two stochastic matrices as
$\bm{P}_{1}$ $=\left(\bm{p}_{1},\ldots,\bm{p}_{n}\right)$ (29)
$\bm{P}_{2}$ $=\left(\bm{q}_{1},\ldots,\bm{q}_{n}\right)$
Each $\bm{p}_{i}$ and $\bm{q}_{i}$ are probability vectors in their own right, with dimension $n$, and each matrix are therefore elements of the product space $\Delta\times\cdots\times\Delta$. This means we can construct the induced minimum length geodesic curve distance from each underlying simplex space. In other words,
$d\left(\left(\bm{p}_{1},\ldots,\bm{p}_{n}\right),\left(\bm{q}_{1},\ldots,\bm{q}_{n}\right)\right)=\sqrt{d^{2}(\bm{p}_{1},\bm{q}_{1})+\ldots+d^{2}(\bm{p}_{n},\bm{q}_{n})}$ (30)
As the underlying simplex spaces can be equipped with the Bhattacharyya angle, we can write the induced distance as
$d\left(\left(\bm{p}_{1},\ldots,\bm{p}_{n}\right),\left(\bm{q}_{1},\ldots,\bm{q}_{n}\right)\right)=2\sqrt{\sum_{i=1}^{n}\arccos^{2}\mathrm{BC}(\bm{p}_{i},\bm{q}_{i})}$ (31)
Noting $\bm{p}_{i}=(p_{i1},\ldots,p_{in})$, $\bm{q}_{1}=(q_{i1},\ldots,q_{in})$, where $p_{ij}=P_{1}(x_{i},x_{j})$ and $q_{ij}=P_{2}(x_{i},x_{j})$, it is easy to see the Bhattacharyya coefficient expressions can be written as $\mathrm{BC}(\bm{p}_{i},\bm{q}_{i})=\sum_{j=1}^{n}\sqrt{p_{ij}q_{ij}}$ which allows us to write
$d\left(\left(\bm{p}_{1},\ldots,\bm{p}_{n}\right),\left(\bm{q}_{1},\ldots,\bm{q}_{n}\right)\right)=2\sqrt{\sum_{x_{i}\in X}\arccos^{2}\sum_{x_{j}\in X}\sqrt{P_{1}(x_{i},x_{j})P_{2}(x_{i},x_{j})}}$ (32)